CHAPTER 11: CHAPTER END REVIEW A 111 LO Best Sleep Company reported the following results from last years operations: SalesWhat is last year's Margin?

What is last year's turnover?

What is last year's return on investment?

What is the margin related to this year's investment?

Show work pleaseeee
Answer #1

What is last year's Margin?

Margin=net income/sales=270,000/1,500,000=18%

What is last year's turnover?

Turnover=Sales=1,500,000

What is last year's return on investment?

Return on investment=net income/investment=270000/1000000=27%

What is the margin related to this year's investment?

Margin=Contribution margin-fixed expenses=70% of sales-144,000=70% of 240,000-144,000=168,000-144,000=24,000